# Ajay Muthukkumar6, Arvind Rajan6, Jaya Kolluri2, Amelia Fiske7

### v1

@@ 5 Conclusion and Future works
5.1 Limitations
This study has several limitations that may affect the generalizability and accuracy of the findings: 1. Lack of NER Tagger: Without integrating NER taggers, there is a risk of misclassifying terms or missing context, which could compromise the precision of keyword identification and associations. However, we were limited by the the computational requirements of NER tagging over the entire $ThePile$ dataset. 2. Selection of Diseases: The chosen diseases and keywords are based on normalized concepts and standard disease classification terms. This selection, though extensive, does not encompass the entire spectrum of medical knowledge, potentially skewing the representation of disease demographics. 3. Subgroup Selection: To demonstrate variation across subgroups, terms used in the CDC national and US surveys were used as grouping terms for quantifying subgroup robustness. In practicality, these are surface-level attributes that are overly simplistic. Furthermore, these terms can also be associated with negative stereotyping if used to polarize. Our aim is to demonstrate variation using common terms, however we hope that future work will build on our approach to go deeper into the nuances of subgroup robustness which should inherently be locally designed and locally governed. 4. Real-World Data Constraints: The datasets used to determine real-world disease prevalence are limited by their availability, completeness, and collection biases. This may hinder the assessment of the broader impacts of findings. 5. Template Sensitivity: The model’s output sensitivity to semantic nuances in template design means the set number of templates may not capture all linguistic or contextual variations influencing model logits and bias assessment. 6. API access model evaluation: Because most API providers do not provide logit access to models nor model weights, these models cannot be evaluated the same way as we evaluated open-weight models. Therefore, we did not include any API-only access model research. 7. Assessing knowledge in pretraining data and models: Other ways to assess knowledge represented in pretraining data and model representations include investigating direct statements about prevalence in the pretraining data and querying prevalence rates from the model. However, we were interested in the more general question of how general distributions in pretraining data contribute to model biases concerning medical reasoning, more broadly, beyond factoid knowledge.
5.2 Future Work
Future research will prioritise 1. Development of Comprehensive Datasets: Efforts will be made to create and employ datasets that furnish more accurate and exhaustive real-world prevalence data for diseases, especially those poorly represented in existing datasets. 2. Impact on Clinical Decision-Making: We plan to investigate the effects of model biases on downstream tasks and clinical decision-making to improve model training and evaluation to mitigate negative impacts.
5.3 Conclusion
@@
5.4 Ethics
Our study at the intersection of AI, bias, and healthcare raises several ethical issues. Our race/ethnicity and gender categorizations are necessarily simplistic and imperfect. For example, we cannot discriminate between biological sex and gender identity. More work will be needed to investigate the full range of identities, including intersectionality. As interrogation of pretraining datasets becomes more refined, it could inadvertently reveal identifying information and infringe on peoples’ privacy, which is particularly important in the sensitive medical domain. Finally, developing these benchmarks and datasets in resource-poor languages is challenging and merits special priories to ensure safe model development for all populations.
5.5 Contribution
Shan Chen and Jack Gallifant contributed equally to this work. They developed the main theoretical framework, performed the experiments, and led the writing of the manuscript.
Mingye Gao and Pedro Moreira contributed equally, focusing on refining the framework, visualizations, and interpreting the results.
Nikolaj Munch was instrumental in experimental design and data collection.
Ajay Muthukkumar, Arvind Rajan, and Jaya Kolluri were key in the disease prevalence data collection, with Ajay taking a leadership role.
Amelia Fiske ensured that the study adhered to the highest ethical standards and was responsible for overseeing the ethical compliance of the research methodology.
Janna Hastings, Hugo Aerts, Brian Anthony, and Leo Celi provided supervision and were involved in the strategic direction of the research and securing funding.
William G. La Cava provided supervision, mentorship of junior team member, and was instrumental in the strategic direction of the research.
Danielle S Bitterman, the corresponding author, oversaw the entire project, coordinated the interdisciplinary team, and secured funding. She mentored junior team members and ensured the final approval of the version to be published.
5.6 Acknowledgement
@@
The authors also acknowledge financial support from the Woods Foundation (DB, SC, HA) NIH (NIH-USA U54CA274516-01A1 (SC, HA, DB), NIH-USA U24CA194354 (HA), NIH-USA U01CA190234 (HA), NIH-USA U01CA209414 (HA), and NIH-USA R35CA22052 (HA), NIH-USA U54 TW012043-01 (JG, LAC), NIH-USA OT2OD032701 (JG, LAC), NIH-USA R01EB017205 (LAC), DS-I Africa U54 TW012043-01 (LAC), Bridge2AI OT2OD032701 (LAC), NSF ITEST 2148451 (LAC) and the European Union - European Research Council (HA: 866504)
NeurIPS Paper Checklist
  1. 1.
    Claims
  3. Answer: [Yes]
  5. 2.
    Limitations
  7. Answer: [Yes]
  9. 3.
    Theory Assumptions and Proofs
  11. Answer: [N/A] .
  13. 4.
    Experimental Result Reproducibility
  15. Answer: [Yes]
  17. 5.
    Open access to data and code
  19. Answer: [Yes]
  21. 6.
    Experimental Setting/Details
  23. Answer:[Yes]
  25. 7.
    Experiment Statistical Significance
  27. Answer: [N/A]
  29. 8.
    Experiments Compute Resources
  31. Answer: [Yes]
  33. 9.
    Code Of Ethics
  35. Answer: [Yes]
  37. 10.
    Broader Impacts
  39. Answer: [Yes]
  41. 11.
    Safeguards
  43. Answer: [N/A]
  45. 12.
    Licenses for existing assets
  47. Answer: [Yes]
  49. 13.
    New Assets
  51. Answer: [Yes]
  53. 14.
    Crowdsourcing and Research with Human Subjects
  55. Answer: [Yes]
  57. 15.
  59. Answer: [N/A]
@@ A.1 Data Collection for Real-World Prevalence
A.2 Comparison of Raw Pile Counts and Real-World Prevalence
@@
A.3 Comparison of Raw Pile Counts, Real-World Prevalence and Llama3 70b rankings
@@ B.2.2 Kendall’s Tau Analysis by Disease Mention Quartiles
The Kendall Tau analysis, as presented in Figures 10 and 11, underscored the lack of significant variation in the correlation of logits to disease co-occurrence counts across different quartiles for both race/ethnicity and gender. This observation suggests that the frequency of disease co-occurrence counts within the Pile dataset did not necessarily enhance the model’s predictive alignment with real-world demographic distributions of disease prevalence. We only show Pythia here because Mamba series of models are not trained on top of deduplicated $Pile$.
@@ C.2.1 Mistral Model Analysis
English Mistral-7b Base 0 51 0 0 0 38 N/A -0.11
Mistral-Instruct RLHF 0 62 0 0 0 27 0.57 -0.05
Mistral-sft 7b SFT 0 2 0 0 0 87 0.40 -0.23
Zephyr-7b DPO 0 6 0 0 0 83 0.40 -0.23
Bio-mistral 7b Biomed 0 74 0 0 0 15 0.37 -0.04
Chinese Mistral-7b Base 0 1 0 0 0 88 N/A -0.13
Mistral-Instruct RLHF 0 89 0 0 0 0 -0.20 0.00
Mistral-sft 7b SFT 0 0 0 0 0 89 1.0 -0.13
Zephyr-7b DPO 0 0 0 0 0 89 1.0 -0.13
Bio-mistral 7b Biomed 0 30 0 0 0 59 0.53 -0.12
Spanish Mistral-7b Base 0 51 0 0 0 38 N/A 0.24
Mistral-Instruct RLHF 0 82 0 0 0 7 0.40 0.25
Mistral-sft 7b SFT 0 47 0 0 0 42 0.60 0.11
Zephyr-7b DPO 0 57 0 0 0 32 0.40 0.17
Bio-mistral 7b Biomed 0 89 0 0 0 0 0.47 0.32
French Mistral-7b Base 0 0 0 0 0 89 N/A -0.21
Mistral-Instruct RLHF 0 0 0 0 0 89 0.99 -0.23
Mistral-sft 7b SFT 0 0 0 0 0 89 0.99 -0.23
Zephyr-7b DPO 0 0 0 0 0 89 0.99 -0.23
Bio-mistral 7b Biomed 0 0 0 0 0 89 0.99 -0.23
@@
English Mistral-7b Base 47 42 0 N/A -0.33
Mistral-Instruct RLHF 0 89 0 0.65 -0.07
Mistral-sft 7b SFT 51 38 0 0.85 -0.07
Zephyr-7b DPO 35 54 0 0.84 -0.33
Bio-mistral 7b Biomed 20 69 0 0.71 -0.33
Chinese Mistral-7b Base 0 89 0 N/A -0.07
Mistral-Instruct RLHF 0 89 0 0.95 -0.07
Mistral-sft 7b SFT 0 89 0 0.95 -0.07
Zephyr-7b DPO 0 89 0 0.95 -0.07
Bio-mistral 7b Biomed 0 89 0 0.93 -0.07
Spanish Mistral-7b Base 89 0 0 N/A 0.07
Mistral-Instruct RLHF 61 0 28 0.80 0.07
Mistral-sft 7b SFT 85 0 4 0.97 0.07
Zephyr-7b DPO 84 0 5 0.96 0.07
Bio-mistral 7b Biomed 5 0 84 0.37 0.07
French Mistral-7b Base 89 0 0 N/A 0.07
Mistral-Instruct RLHF 89 0 0 1.0 0.07
Mistral-sft 7b SFT 89 0 0 1.0 0.07
Zephyr-7b DPO 89 0 0 1.0 0.07
Bio-mistral 7b Biomed 89 0 0 1.0 0.07
@@ C.2.2 Qwen Model Analysis
Similarly, the Qwen series models demonstrated how different configurations influence demographic trends in model outputs. Figure 15, Figure 16 and Figure 17 highlight these trends, providing insight into the effectiveness of the model’s alignment and training data in reflecting diverse demographic attributes. Table 6 shows the model’s top demographic choices, drift from base models, and Kendall’s tau score compared to real-world prevalence.
@@
Language Model Name Alignment A B H I PI W $\delta$ $\delta\uparrow$ $\tau\uparrow$
English Qwen1.5-7b Base 78 0 0 11 0 0 N/A 0.00
Qwen1.5-7b chat RLHF 5 0 0 84 0 0 0.17 -0.19
Qwen1.5-72b Base 86 0 0 3 0 0 0.49 0.05
Qwen1.5-72b chat RLHF 85 4 0 0 0 0 0.45 0.04
Chinese Qwen1.5-7b Base 89 0 0 0 0 0 N/A -0.03
Qwen1.5-7b chat RLHF 79 10 0 0 0 0 0.83 0.04
Qwen1.5-72b Base 89 0 0 0 0 0 -0.16 0.07
Qwen1.5-72b chat RLHF 86 0 0 0 0 3 0.00 0.01
Spanish Qwen1.5-7b Base 0 52 0 0 0 37 N/A -0.16
Qwen1.5-7b chat RLHF 0 56 0 0 0 33 0.60 -0.13
Qwen1.5-72b Base 0 78 0 0 0 11 0.40 -0.07
Qwen1.5-72b chat RLHF 0 26 0 0 0 63 0.60 -0.21
French Qwen1.5-7b Base 0 89 0 0 0 0 N/A 0.05
Qwen1.5-7b chat RLHF 0 89 0 0 0 0 0.91 0.01
Qwen1.5-72b Base 0 89 0 0 0 0 0.99 0.04
Qwen1.5-72b chat RLHF 0 88 0 0 0 1 0.92 0.03
@@
English Qwen1.5-7b Base 3 86 0 N/A -0.07
Qwen1.5-7b chat RLHF 0 89 0 0.98 -0.07
Qwen1.5-72b Base 25 64 0 0.79 -0.07
Qwen1.5-72b chat RLHF 28 61 0 0.78 -0.20
Chinese Qwen1.5-7b Base 89 0 0 N/A 0.07
Qwen1.5-7b chat RLHF 89 0 0 1.0 0.07
Qwen1.5-72b Base 89 0 0 1.0 0.07
Qwen1.5-72b chat RLHF 89 0 0 1.0 0.07
Spanish Qwen1.5-7b Base 85 0 4 N/A 0.07
Qwen1.5-7b chat RLHF 86 0 3 0.96 0.07
Qwen1.5-72b Base 89 0 0 0.97 0.07
Qwen1.5-72b chat RLHF 47 0 42 0.69 0.07
French Qwen1.5-7b Base 89 0 0 N/A 0.07
Qwen1.5-7b chat RLHF 89 0 0 1.0 0.07
Qwen1.5-72b Base 89 0 0 -0.62 0.07
Qwen1.5-72b chat RLHF 89 0 0 0.6 0.07
@@
C.2.3 LLama3 Model Analysis

### v2

@@ 5 Conclusion and Future works
Limitations
This study has limitations that should be considered when interpreting the findings:
1. Lack of NER Tagger: Without integrating NER taggers, there is a risk of misclassifying terms or missing context. However, we were limited by the the computational requirements of NER tagging over the entire $ThePile$ dataset.
2. Selection of Diseases: The chosen diseases and keywords are based on normalized concepts and standard disease classification terms. This selection, though extensive, does not encompass the entire spectrum of medical knowledge which could skew findings.
3. Subgroup Selection: To demonstrate variation across subgroups, terms used in the CDC national and US surveys were used as grouping terms for quantifying subgroup robustness. These terms are surface-level attributes that are overly simplistic and can also be associated with negative stereotyping if used to polarize. Our aim is to demonstrate variation using common terms, however we hope that future work will build on our approach to go deeper into the nuances of subgroup robustness which should inherently be locally designed and locally governed.
4. Real-World Data Constraints: The datasets used to determine real-world disease prevalence are limited by their availability, completeness, and collection biases. This may hinder the assessment of the broader impacts of findings.
5. Template Sensitivity: The model’s output sensitivity to semantic nuances in template design means the set number of templates may not capture all linguistic or contextual variations influencing model logits and bias assessment.
6. API access model evaluation: Because most API providers do not provide logit access to models nor model weights, these models cannot be evaluated the same way as we evaluated open-weight models. Therefore, we did not include any API-only access model research.
7. Assessing knowledge in pretraining data and models: Other ways to assess knowledge represented in pretraining data and model representations include investigating direct statements about prevalence in the pretraining data and querying prevalence rates from the model. However, we were interested in the more general question of how general distributions in pretraining data contribute to model biases concerning medical reasoning, more broadly, beyond factoid knowledge.
Future Work
Future research will prioritise:
1. Development of Comprehensive Datasets: Efforts will be made to create and employ datasets that furnish more accurate and exhaustive real-world prevalence data for diseases, especially those poorly represented in existing datasets.
2. Impact on Clinical Decision-Making: We plan to investigate the effects of model biases on downstream tasks and clinical decision-making to improve model training and evaluation to mitigate negative impacts.
Conclusion
@@
Acknowledgement
@@
The authors also acknowledge financial support from the Woods Foundation (DB, SC, HA) NIH (NIH-USA U54CA274516-01A1 (SC, HA, DB), NIH-USA U24CA194354 (HA), NIH-USA U01CA190234 (HA), NIH-USA U01CA209414 (HA), and NIH-USA R35CA22052 (HA), NIH-USA U54 TW012043-01 (JG, LAC), NIH-USA OT2OD032701 (JG, LAC), NIH-USA R01EB017205 (LAC), NIH-USA R01LM014300 (WGL), DS-I Africa U54 TW012043-01 (LAC), Bridge2AI OT2OD032701 (LAC), NSF ITEST 2148451 (LAC) and the European Union - European Research Council (HA: 866504)
@@ A.1 Data Collection for Real-World Prevalence
A.2 Definition of Kendal Tau and usage in ranking comparison
We computed the Kendall’s tau scores to compare the rank order of diseases based on:
Model Logit Ranks versus Pile Demographic Subgroup Co-occurrence Ranks: This comparison assessed how model predictions align with the observed co-occurrence frequencies of disease co-occurrence counts within specific demographic subgroups in the training data.
Model Logit Ranks versus Real-World Gold Subset Ranks: This analysis examined the alignment of model outputs with disease rankings derived from a curated real-world dataset, providing insights into the model’s ability to mirror actual disease prevalence across different demographic subgroups.
A.3 Comparison of Raw Pile Counts, Real-World Prevalence
@@
Comparison of Raw Pile Counts, Real-World Prevalence and Llama3 70b rankings
@@ B.2.2 Kendall’s Tau Analysis by Disease Mention Quartiles
The Kendall Tau analysis, as presented in Figures 9 and 10, underscored the lack of significant variation in the correlation of logits to disease co-occurrence counts across different quartiles for both race/ethnicity and gender. This observation suggests that the frequency of disease co-occurrence counts within the Pile dataset did not necessarily enhance the model’s predictive alignment with real-world demographic distributions of disease prevalence. We only show Pythia here because Mamba series of models are not trained on top of deduplicated $Pile$.
@@ C.2.1 Mistral Model Analysis
English Mistral-7b Base 0 51 0 0 0 38 N/A -0.20
Mistral-Instruct RLHF 0 62 0 0 0 27 0.91 -0.12
Mistral-sft 7b SFT 0 2 0 0 0 87 0.92 -0.13
Zephyr-7b DPO 0 6 0 0 0 83 0.92 -0.13
Bio-mistral 7b Biomed 0 74 0 0 0 15 0.88 -0.05
Chinese Mistral-7b Base 0 1 0 0 0 88 N/A 0.02
Mistral-Instruct RLHF 0 89 0 0 0 0 0.82 0.04
Mistral-sft 7b SFT 0 0 0 0 0 89 1.0 0.02
Zephyr-7b DPO 0 0 0 0 0 89 1.0 0.02
Bio-mistral 7b Biomed 0 30 0 0 0 59 0.94 0.03
Spanish Mistral-7b Base 0 51 0 0 0 38 N/A 0.18
Mistral-Instruct RLHF 0 82 0 0 0 7 0.92 0.22
Mistral-sft 7b SFT 0 47 0 0 0 42 0.95 0.21
Zephyr-7b DPO 0 57 0 0 0 32 0.92 0.22
Bio-mistral 7b Biomed 0 89 0 0 0 0 0.93 0.21
French Mistral-7b Base 0 0 0 0 0 89 N/A -0.12
Mistral-Instruct RLHF 0 0 0 0 0 89 0.99 -0.13
Mistral-sft 7b SFT 0 0 0 0 0 89 0.99 -0.13
Zephyr-7b DPO 0 0 0 0 0 89 0.99 -0.13
Bio-mistral 7b Biomed 0 0 0 0 0 89 0.99 -0.13
@@
English Mistral-7b Base 47 42 0 N/A -0.47
Mistral-Instruct RLHF 0 89 0 0.65 -0.20
Mistral-sft 7b SFT 51 38 0 0.85 -0.20
Zephyr-7b DPO 35 54 0 0.84 -0.47
Bio-mistral 7b Biomed 20 69 0 0.71 -0.47
Chinese Mistral-7b Base 0 89 0 N/A -0.20
Mistral-Instruct RLHF 0 89 0 0.95 -0.20
Mistral-sft 7b SFT 0 89 0 0.95 -0.20
Zephyr-7b DPO 0 89 0 0.95 -0.20
Bio-mistral 7b Biomed 0 89 0 0.93 -0.20
Spanish Mistral-7b Base 89 0 0 N/A 0.20
Mistral-Instruct RLHF 61 0 28 0.80 0.20
Mistral-sft 7b SFT 85 0 4 0.97 0.20
Zephyr-7b DPO 84 0 5 0.96 0.20
Bio-mistral 7b Biomed 5 0 84 0.37 0.20
French Mistral-7b Base 89 0 0 N/A 0.20
Mistral-Instruct RLHF 89 0 0 1.0 0.20
Mistral-sft 7b SFT 89 0 0 1.0 0.20
Zephyr-7b DPO 89 0 0 1.0 0.20
Bio-mistral 7b Biomed 89 0 0 1.0 0.20
@@
$\tau:\{-1,1\}$ Kendall Tau of model’s prevalence representation vs real-world prevalence
@@ C.2.2 Qwen Model Analysis
Similarly, the Qwen series models demonstrated how different configurations influence demographic trends in model outputs. Figure 14, Figure 15 and Figure 16 highlight these trends, providing insight into the effectiveness of the model’s alignment and training data in reflecting diverse demographic attributes. Table 5(b) shows the model’s top demographic choices, drift from base models, and Kendall’s tau score compared to real-world prevalence.
@@
Language Model Name Alignment A B H I PI W $\delta$uparrow $\tau\uparrow$
English Qwen1.5-7b Base 78 0 0 11 0 0 N/A -0.02
Qwen1.5-7b chat RLHF 5 0 0 84 0 0 0.88 0.11
Qwen1.5-72b Base 86 0 0 3 0 0 0.89 -0.10
Qwen1.5-72b chat RLHF 85 4 0 0 0 0 0.83 -0.17
Chinese Qwen1.5-7b Base 89 0 0 0 0 0 N/A -0.42
Qwen1.5-7b chat RLHF 79 10 0 0 0 0 0.96 -0.44
Qwen1.5-72b Base 89 0 0 0 0 0 0.76 -0.28
Qwen1.5-72b chat RLHF 86 0 0 0 0 3 0.85 -0.41
Spanish Qwen1.5-7b Base 0 52 0 0 0 37 N/A -0.14
Qwen1.5-7b chat RLHF 0 56 0 0 0 33 0.95 -0.17
Qwen1.5-72b Base 0 78 0 0 0 11 0.92 -0.16
Qwen1.5-72b chat RLHF 0 26 0 0 0 63 0.95 -0.14
French Qwen1.5-7b Base 0 89 0 0 0 0 N/A 0.00
Qwen1.5-7b chat RLHF 0 89 0 0 0 0 0.94 -0.05
Qwen1.5-72b Base 0 89 0 0 0 0 0.99 0.00
Qwen1.5-72b chat RLHF 0 88 0 0 0 1 0.98 0.02
@@
English Qwen1.5-7b Base 3 86 0 N/A -0.20
Qwen1.5-7b chat RLHF 0 89 0 0.98 -0.20
Qwen1.5-72b Base 25 64 0 0.79 -0.20
Qwen1.5-72b chat RLHF 28 61 0 0.78 -0.33
Chinese Qwen1.5-7b Base 89 0 0 N/A 0.20
Qwen1.5-7b chat RLHF 89 0 0 1.0 0.20
Qwen1.5-72b Base 89 0 0 1.0 0.20
Qwen1.5-72b chat RLHF 89 0 0 1.0 0.20
Spanish Qwen1.5-7b Base 85 0 4 N/A 0.20
Qwen1.5-7b chat RLHF 86 0 3 0.96 0.20
Qwen1.5-72b Base 89 0 0 0.97 0.20
Qwen1.5-72b chat RLHF 47 0 42 0.69 0.20
French Qwen1.5-7b Base 89 0 0 N/A 0.20
Qwen1.5-7b chat RLHF 89 0 0 1.0 0.20
Qwen1.5-72b Base 89 0 0 0.46 0.20
Qwen1.5-72b chat RLHF 89 0 0 0.87 0.20
@@
$\tau:\{-1,1\}$ Kendall Tau of model’s prevalence representation vs real-world prevalence
C.2.3 Llama2 Model Analysis
The Llama-2 series models demonstrated how different configurations influence demographic trends in model outputs. Figure 17, Figure 18 and Figure 19 highlight these trends, providing insight into the effectiveness of the model’s alignment and training data in reflecting diverse demographic attributes. Table 6(b) shows the model’s top demographic choices, drift from base models, and Kendall’s tau score compared to real-world prevalence.
Figure 18: Bottom ranked gender (top) and race/ethnicity (bottom) subgroups across each of the 89 diseases using the Llama series across 4 languages. en, English; zh, Mandarin; es, Spanish; fr, French
Figure 19: Second bottom ranked gender (top) and race/ethnicity (bottom) subgroups across each of the 89 diseases using the Llama series across 4 languages. en, English; zh, Mandarin; es, Spanish; fr, French
Language Model Name Alignment A B H I PI W $\delta\uparrow$ $\tau\uparrow$
English Llama2-70b Base 0 70 0 0 0 19 N/A -0.17
Llama2-70b chat RLHF 0 86 0 0 0 3 0.91 -0.17
Tulu2-70b SFT 0 65 0 0 0 24 0.98 -0.14
Tulu2-70b-dpo DPO 1 54 0 0 0 34 0.98 -0.14
Meditron-70b Biomed 0 60 0 0 0 29 0.95 -0.14
Chinese Llama2-70b Base 0 0 0 0 0 89 N/A 0.02
Llama2-70b chat RLHF 0 0 0 0 0 89 0.99 0.03
Tulu2-70b SFT 0 0 0 0 0 89 0.99 0.00
Tulu2-70b-dpo DPO 0 0 0 0 0 89 0.98 0.00
Meditron-70b Biomed 0 0 0 0 0 89 1.0 0.02
Spanish Llama2-70b Base 0 89 0 0 0 0 N/A 0.00
Llama2-70b chat RLHF 0 89 0 0 0 0 1.0 0.00
Tulu2-70b SFT 0 89 0 0 0 0 1.0 0.00
Tulu2-70b-dpo DPO 0 89 0 0 0 0 1.0 0.00
Meditron-70b Biomed 0 89 0 0 0 0 0.99 -0.01
French Llama2-70b Base 0 0 0 0 0 89 N/A 0.02
Llama2-70b chat RLHF 0 0 0 0 0 89 0.85 0.22
Tulu2-70b SFT 0 0 0 0 0 89 0.97 0.06
Tulu2-70b-dpo DPO 0 0 0 0 0 89 0.97 0.06
Meditron-70b Biomed 0 0 0 0 0 89 0.88 -0.13
(a) Demographic distribution by race and model alignment
English Llama2-70b Base 36 53 0 N/A -0.60
Llama2-70b chat RLHF 19 70 0 0.77 0.33
Tulu2-70b SFT 22 67 0 0.85 -0.33
Tulu2-70b-dpo DPO 10 79 0 0.79 -0.20
Meditron-70b Biomed 16 73 0 0.79 -0.20
Chinese Llama2-70b Base 88 1 0 N/A 0.07
Llama2-70b chat RLHF 1 0 88 -0.53 -0.60
Tulu2-70b SFT 53 34 2 0.67 -0.73
Tulu2-70b-dpo DPO 49 39 1 0.66 -0.73
Meditron-70b Biomed 75 14 0 0.9 -0.47
Spanish Llama2-70b Base 44 0 45 N/A 0.20
Llama2-70b chat RLHF 83 0 6 0.68 0.20
Tulu2-70b SFT 85 0 4 0.69 0.20
Tulu2-70b-dpo DPO 87 0 2 0.68 0.20
Meditron-70b Biomed 89 0 0 0.66 0.20
French Llama2-70b Base 0 89 0 N/A -0.20
Llama2-70b chat RLHF 9 80 0 0.93 -0.47
Tulu2-70b SFT 0 89 0 1.0 -0.20
Tulu2-70b-dpo DPO 0 89 0 1.0 -0.20
Meditron-70b Biomed 0 89 0 1.0 -0.20
(b) Demographic distribution by gender and model alignment
$\tau:\{-1,1\}$ Kendall Tau of model’s prevalence representation vs real-world prevalence
C.2.4 Llama3 Model Analysis